OpenStudy (anonymous):

find the zeros of 4x-(1/(2x^2))

jhonyy9 (jhonyy9):

x not can being zero

jhonyy9 (jhonyy9):

8x3 -1=0 8x3=1 x3=1/8 x=cuberoot(1/8) = 1/2

OpenStudy (anonymous):

where did u get the 8?

OpenStudy (anonymous):

oh nvm ty

jhonyy9 (jhonyy9):

ok ?